# AugGen: Augmenting Task-Based Learning in Professional Creative Software with LLM-Generated Scaffolded UIs

### v1

@@ 7.2. Limitations and Future Directions
Our work has some limitations that we discuss below. First, our interface uses user-selectable complexity levels for progressive tool disclosure. This is just one point on the spectrum from user-driven to system-driven control. Future research could explore system-driven interface adaptation based on performance metrics or user modeling, as well as mixed-initiative approaches where the system suggests level changes or adapts specific interface elements contextually. Second, some expert participants raised concerns about potential over-reliance on the scaffolded interface, which might limit learning of the full software. To address this, future design could include cues like UI minimaps or short animations showing where tools come from. The scaffolded interface could also fade as users gain skill, encouraging further transition to the native UI. Third, expert feedback suggested that they desired UI customization for personalized workflows. This encourages the exploration of combining on-demand and scaffolded UI generation. Future research could use our technical pipeline, which allows interface iteration with human control, to let users customize their interfaces and evaluate such dynamic user experience. Lastly, our technical implementation demonstrated an example of building scaffolded interfaces in a software tool for 3D modeling and animation tasks. As introduced in Section 3, our approach can generate UIs for broader tasks and software that offers manuals and APIs. Future research could test the implementation of scaffolded interfaces in other applications and domains to explore broader use and impact.
@@ 8. Conclusion
Steep learning curves in professional creative software due to complex interfaces significantly hinder user learning and skill development. In this paper, we introduced AugGen, a method for generating scaffolded interfaces in professional creative software to support task-based learning. The scaffolded interfaces present task-relevant tools, manage complexity through adjustable levels, organize the UI around workflow stages with domain concepts integrated, and connect users to native software interactions to support skill transfer. Our evaluation involving both beginners and experts showed that the scaffolded interfaces significantly improved task performance for both user groups and supported learning during task execution for beginners. These findings demonstrated the effectiveness of combining task-awareness, workflow-based organization, concept integration, and progressive feature disclosure in interface design. Based on a discussion of the study results, we provided insights for future research on user empowerment through scaffolded interfaces for software learners, educators, and end users.
@@ A.1. Workflow Analysis and Concept Identification
You are an expert task workflow analyst specializing in task decomposition for a user task. Your objective is to analyze the user task and break it down into its primary workflow stages. TASK TO ANALYZE: {USER TASK DESCRIPTION} METHOD: Decompose the task into logical stages that represent a typical, efficient workflow a user would follow within {SOFTWARE}. Focus on the major distinct phases, moving from initial setup/preparation towards finalization/review. Aim for a reasonable number of stages that capture the core workflow without excessive granularity. Also, extract domain concepts relevant to each workflow stage. For each domain concepts, provide a concise, straightforward explanation for a human to easily understand. OUTPUT REQUIREMENTS: Provide the output as a numbered list. Each list item must represent a distinct workflow stage and include: A concise Stage Name. A brief, one-sentence Stage Description clarifying the purpose of the stage. Domain Concepts related to the stage. A concise, straightforward Concept Explanation explaining the concept. The output must be structured following the EXAMPLE OUTPUT FORMAT. EXAMPLE OUTPUT FORMAT: Stage Name 1: Stage Description, Domain Concepts, Concept Explanation. Stage Name 2: Stage Description, Domain Concepts, Concept Explanation. Stage Name n: Stage Description, Domain Concepts, Concept Explanation.
@@ A.2. Tool Selection and Complexity Assessment
You are an expert user of {SOFTWARE}. Your task is to identify relevant tools in {SOFTWARE} to perform each workflow stage of {WORKFLOW STAGES}. WORKFLOW STAGES: {STAGE NAME}: {STAGE DESCRIPTION}, {DOMAIN CONCEPTS}, {CONCEPT EXPLANATION} METHOD: Identify the keyboard operators, mouse operations, and UI interactions needed to perform each workflow stage. The identified tools must be common and essential for the workflow stage. Prioritize stable and frequently used tools. In addition, you need to assign a difficulty level for each identified tool by referring to {SOFTWARE MANUALS}. The difficulty level should be basic, intermediate, or advanced for a human user. OUTPUT REQUIREMENTS: For each stage, output a bulleted list of selected tools. Each item in the list must include: The Stage Name from the {WORKFLOW STAGES}. The Selected Tools required to perform the task in the stage. A Rationale that briefly and clearly explain why the selected tools are essential and necessary. A Complexity Level of Basic, Intermediate, or Advanced for the selected tools. EXAMPLE OUTPUT FORMAT: Stage Name 1: Selected Tools, Rationale, Complexity Level. Stage Name 2: Selected Tools, Rationale, Complexity Level. Stage Name n: Selected Tools, Rationale, Complexity Level.
@@ A.3. UI Code Generation and Tool Labeling
You are a coding expert for {SOFTWARE}. Your task is to generate the code to implement {SELECTED TOOLS} based on {SOFTWARE} functionality and generate UI code to display {SELECTED TOOLS} in a user interface for user interaction. SELECTED TOOLS: {STAGE NAME}: {SELECTED TOOLS}, {RATIONALE}, {COMPLEXITY LEVEL}. METHOD: For each selected tool in {SELECTED TOOLS}, you need to translate the tool operation into the corresponding code execution by referring to {SOFTWARE APIS}. You must make sure the code can achieve the identical result using the tool. The code you generate should include all the required parameters and properties for the code to be correctly executed. Next, you need to convert the code into interactive UI controls that can be used by a human. You can use buttons, dropdown menus, radio buttons, tabs, text field, switch toggles, or other UI controls that are supported by {SOFTWARE} and suitable for each tool. You must refer to {SOFTWARE APIS} to make sure the UI control can be implemented. To organize the code for UI controls, you should refer to the {EXAMPLE CODE} to group the code for each UI control under their {WORKFLOW STAGE} and {COMPLEXITY LEVEL}. In addition, you should implement a mouse-over tooltip for each UI control with their relevant {DOMAIN CONCEPT} and {CONCEPT DESCRIPTION}. EXAMPLE CODE WITH UI LAYOUT STRUCTURE {EXAMPLE CODE} OUTPUT REQUIREMENTS: Your output is a code script. The script should be standalone with all necessary imports, function definitions, and main function for it to be executable. You should refer to {EXAMPLE CODE} for the code script generation.
@@ B.1. Walk Cycle Scaffolded UI
Figure 11. Walk cycle animation using the scaffolded UI.:
@@ B.3. Questionnaires
Table 1 summarizes the questions from NASA-TLX to evaluate task load. The custom questionnaires for task performance, concept learning, and user interface experience are presented in Tables 2, 3, and 4.
@@ B.4. Open-Ended Questions
Appendix C User Study 2 Participant Demographic Information
Table 5 presents the demographic information of Study 2 participants.
Table 5. Study 2 participant demographic information.
S2-P1 2 SketchUp
S2-P2 5 Maya, ZBrush, MotionBuilder
S2-P3 3 Maya
S2-P4 3 3ds Max, Mixamo, Unity, Unreal Engine
S2-P5 6 Unity, Maya
S2-P6 4 Tinkercad, 3ds Max, Maya
S2-P7 1 FreeCAD, Tinkercad
S2-P8 5 Substance 3D, After Effects
Appendix D User Study 1 ANOVA Results
@@
Appendix E Scaffolded Interface Variations
E.1. UV Unwrapping
@@
E.2. Building Walk Cycle
@@
Figure 12. AugGen variations.:

### v2

@@ 7.3. Limitations and Future Directions
We discuss our work’s limitations here and offer insights for future research. First, our interface uses user-selectable complexity levels for progressive tool disclosure. This is just one point on the spectrum from user-driven to system-driven control. Future research could explore system-driven interface adaptation based on performance metrics or user modeling, as well as mixed-initiative approaches where the system suggests level changes or adapts specific interface elements contextually. Second, expert feedback suggested that they desired UI customization for personalized workflows. This encourages the exploration of combining on-demand and scaffolded UI generation. Future research could use our technical pipeline, which allows interface iteration with human control, to let users customize their interfaces and evaluate such dynamic user experience. Third, we evaluated task-conditioned scaffolded interfaces in Blender for 3D modeling and animation tasks. As introduced in Section 3.2.3 and discussed in Section 7.1, our method can be used to generate scaffolded UIs for broader tasks and in other software tools. Future research could evaluate scaffolded interfaces in other applications and domains to explore broader use and impact. Lastly, we tested GPT-4o as the LLM to analyze user task, select tools, identify domain concepts, and generate UI implementation code. Future work can evaluate our method using more advanced LLMs, which may reason tools and concepts that are more aligned with UI designers’, software educators’, or end users’ requirements with improved reasoning and instruction following capabilities. In addition, LLMs adept at generating code can help to boost the UI implementation reliability, thus enabling easier UI iteration and customization.
@@ 8. Conclusion
Steep learning curves in professional creative software due to complex interfaces significantly hinder user learning and skill development. In this paper, we introduced TaskLens, a method for generating scaffolded interfaces in professional creative software to support task-based learning. The scaffolded interfaces present task-relevant tools, manage complexity through adjustable levels, organize the UI around workflow stages with domain concepts integrated, and connect users to native software interactions to support skill transfer. Our evaluation involving both beginners and experts showed that the scaffolded interfaces significantly improved task performance and efficiency for both user groups, supported concept understanding and learning during task execution for beginners, and demonstrated potential to create personalized UIs to augment productivity and creativity for experts. These findings showed the effectiveness of combining task-awareness, workflow-based organization, concept integration, and progressive feature disclosure in interface design. Based on a discussion of the study results, we provided insights for future research on user empowerment through scaffolded interfaces for software learners, educators, and end users.
Acknowledgements.
We thank Jennifer Jacobs and Yuan-Fang Wang for their helpful suggestions on this project, our user study participants for their time and the valuable insights they contributed to our evaluation, and the anonymous reviewers whose constructive feedback helped us improve this paper.
@@ A.1. Workflow Analysis and Concept Identification
You are an expert task workflow analyst specializing in task decomposition for a user task. Your objective is to analyze the user task and break it down into its primary workflow stages.
TASK TO ANALYZE:
{USER TASK DESCRIPTION}
METHOD:
Decompose the task into logical stages that represent a typical, efficient workflow a user would follow within {SOFTWARE}. Focus on the major distinct phases, moving from initial setup/preparation towards finalization/review. Aim for a reasonable number of stages that capture the core workflow without excessive granularity. Also, extract domain concepts relevant to each workflow stage. For each domain concepts, provide a concise, straightforward explanation for a human to easily understand.
OUTPUT REQUIREMENTS:
Provide the output as a numbered list. Each list item must represent a distinct workflow stage and include: A concise Stage Name. A brief, one-sentence Stage Description clarifying the purpose of the stage. Domain Concepts related to the stage. A concise, straightforward Concept Explanation explaining the concept.
The output must be structured following the EXAMPLE OUTPUT FORMAT.
EXAMPLE OUTPUT FORMAT:
  • Stage Name 1: Stage Description, Domain Concepts, Concept Explanation.
  • Stage Name 2: Stage Description, Domain Concepts, Concept Explanation.
  • Stage Name n: Stage Description, Domain Concepts, Concept Explanation.
@@ A.2. Tool Selection and Complexity Assessment
You are an expert user of {SOFTWARE}. Your task is to identify relevant tools in {SOFTWARE} to perform each workflow stage of {WORKFLOW STAGES}.
WORKFLOW STAGES:
  • {STAGE NAME}: {STAGE DESCRIPTION}, {DOMAIN CONCEPTS}, {CONCEPT EXPLANATION}
METHOD:
Identify the keyboard operators, mouse operations, and UI interactions needed to perform each workflow stage. The identified tools must be common and essential for the workflow stage. Prioritize stable and frequently used tools. In addition, you need to assign a difficulty level for each identified tool by referring to {SOFTWARE MANUALS}. The difficulty level should be basic, intermediate, or advanced for a human user.
OUTPUT REQUIREMENTS:
For each stage, output a bulleted list of selected tools. Each item in the list must include: The Stage Name from the {WORKFLOW STAGES}. The Selected Tools required to perform the task in the stage. A Rationale that briefly and clearly explain why the selected tools are essential and necessary. A Complexity Level of Basic, Intermediate, or Advanced for the selected tools.
EXAMPLE OUTPUT FORMAT:
  • Stage Name 1: Selected Tools, Rationale, Complexity Level.
  • Stage Name 2: Selected Tools, Rationale, Complexity Level.
  • Stage Name n: Selected Tools, Rationale, Complexity Level.
@@ A.3. UI Code Generation and Tool Labeling
You are a coding expert for {SOFTWARE}. Your task is to generate the code to implement {SELECTED TOOLS} based on {SOFTWARE} functionality and generate UI code to display {SELECTED TOOLS} in a user interface for user interaction.
SELECTED TOOLS:
  • {STAGE NAME}: {SELECTED TOOLS}, {RATIONALE}, {COMPLEXITY LEVEL}.
METHOD:
For each selected tool in {SELECTED TOOLS}, you need to translate the tool operation into the corresponding code execution by referring to {SOFTWARE APIS}. You must make sure the code can achieve the identical result using the tool. The code you generate should include all the required parameters and properties for the code to be correctly executed.
Next, you need to convert the code into interactive UI controls that can be used by a human. You can use buttons, dropdown menus, radio buttons, tabs, text field, switch toggles, or other UI controls that are supported by {SOFTWARE} and suitable for each tool. You must refer to {SOFTWARE APIS} to make sure the UI control can be implemented.
To organize the code for UI controls, you should refer to the {EXAMPLE CODE} to group the code for each UI control under their {WORKFLOW STAGE} and {COMPLEXITY LEVEL}. In addition, you should implement a mouse-over tooltip for each UI control with their relevant {DOMAIN CONCEPT} and {CONCEPT DESCRIPTION}.
EXAMPLE CODE WITH UI LAYOUT STRUCTURE
{EXAMPLE CODE}
OUTPUT REQUIREMENTS:
Your output is a code script. The script should be standalone with all necessary imports, function definitions, and main function for it to be executable. You should refer to {EXAMPLE CODE} for the code script generation.
@@ B.1. Walk Cycle Scaffolded UI
Figure 11. Our scaffolded UI generated to perform walk cycle animation.
This figure shows the scaffolded UI for building the walk cycle animation. This scaffolded UI appears as an add-on panel on the right side in Blender’s UI.
@@ B.3. Questionnaires
Table 2 summarizes the questions from NASA-TLX to evaluate task load. The custom questionnaires for task performance, concept learning, and user interface experience are presented in Tables 34, and 5.
@@ B.4. Open-Ended Questions
Appendix C User Study 1 ANOVA Results
@@
Appendix D Scaffolded Interface Variations
D.1. UV Unwrapping
@@
D.2. Building Walk Cycle
@@
Figure 12. Scaffolded UI variations in Blender implemented using TaskLens.
This figure shows the UI variations of the UV unwrapping and walk animation tasks.
Appendix E Scaffolded Interface in Other Software
Figures 13(a) and 13(b) demonstrate the scaffolded interfaces we implemented in Maya to show the possibility of applying our scaffolded UI generation method in other professional software.
(a) UV unwrapping.
(b) Building walk cycle.
Figure 13. Scaffolded UIs in Maya implemented using TaskLens.
This figure shows the scaffolded UIs for the UV unwrapping and walk animation tasks in Maya.